2010 Prices 2011 Prices 2012 Prices
Milk $2 $3 $4
Cookies $2 $3 $2
CPI

A.  For a fixed basket of 5 glasses of milk and 10 cookies, and using the prices in the table, calculate the CPI in each year for this economy. 2010 is your base year.

B.  If the CPI in 2013 was 192 and the CPI in 2014 was 200, what would the inflation rate be? What would the inflation rate mean?

C.  If you earned $10/hour in 2010 (base year) and the CPI in 2015 was 195, what would your 2010 $10/hour wage be worth in 2015 terms?

D.  If the inflation rate in 2015 was -2%, what would be occurring? What would that mean?
Answer #1

A.Total GDP in 2010 = 10*2+5*2=30. CPI= 100

Total GDP in 2011= 3*10+3*5= 45 CPI = 150,(45/30)*100

Total GDP in 2012=2*10+4*5= 40 CPI = 133.3,(40÷30)*100

B.Inflation rate =[ (200/192)*100]-100=4.17%. It means, the av prices grew by 4.17 on an average from 2013 to 2014.

C.10 dollar wage in 2010 would be worth 10*195/100 = 19.5 dollar/hour

D.if inflation in 2015 was 2%, deflation would be occuring, i.e on an average the price level has gone down by 2% from 2014 to 2015 which is indicated by a 2% fall in CPI